# The entropy corrected geometric Brownian motion

### 10 Mar 2024 v1

@@ I Introduction
The geometric Brownian motion (GBM) has long served as a foundational model for capturing stochastic nature of processes characterized by the continuous random fluctuations. In particular, this model has has been argued to be well suited for forecasting diffusion processes [1], population dynamics [2] or most notably stock prices [3]. However, central to GBM is the assumption that logarithm of its solutions result in a normal distribution, a premise that, while convenient for analytical purposes, may not fully encapsulate the complexities inherent in realistic systems [4]. For instance, real data distributions often exhibit characteristics such as non-zero skewness, excess kurtosis or fluctuating volatility, notably deviating from the idealized bell curve [5]. These deviations have significant implications for the accuracy of traditional GBM, by challenging its abilities to properly portray extreme events or interpret underlying dynamics of considered data.
By recognizing these limitations, a motivation arises for the exploration of solutions that go beyond the described constraints of normality. While alternative models, such as the stochastic volatility [6] and the jump-diffusion process [7], have been proposed, they mostly call for the modification or replacement of GBM. Hence, the existing advantages of this framework, in terms of its simplicity, efficiency and continuity, are somewhat hindered or lost. As an alternative, the present study proposes non-violent approach of employing entropy constraints to improve predictive capabilities of GBM without sacrificing its beneficial features. The idea originates from the Shannon’s information theory, where entropy plays pivotal role and stands for the minimum number of logical states needed to communicate some message [8]. This is to say, the well-ordered message is expected to have a lower entropy than a more random one. As a result, it can be argued that entropy provides a means to judge how well GBM predicts future events by measuring the level of uncertainty around forecast data with respect to the original time series. This premise is additionally reinforced when invoking the ability of entropy measure to capture extreme events [9].
To better picture above reasoning let us conceive gedanken experiment that hinges on a dice roll. We begin with a well known fact that probability of getting any side number during a conventional dice roll is equally probable, resulting in a uniform distribution. According to Shannon, the information about such process can be quantified by the entropy given as [8]:
@@ II Methodology
where $S(t)$ is the series value at time $t$, $\mu$ and $\sigma$ are respectively the drift and volatility coefficients, while $dW(t)$ denotes the Wiener process (the Brownian motion). By using Itô’s calculus and assuming log-normal distribution of considered data under GBM, the central equation of $S(t)$ can be obtained [10, 11, 12]:
@@ III Biased dice roll
As a measure of the difference between the reference and generated probability distributions, given by GBM and EC-GBM methods, the Kullback-Leibler ($D_{KL}$) divergence is calculated and shown in Figure 3. The $D_{KL}$ measure is often used in information theory and statistics to quantify the difference between two probability distributions in the following manner [13]:
@@ V Conclusion
In summary, the present research endeavor allows to deepen current understanding of how to effectively discern and harness the wealth of information contained within historical time series. It aligns with the concept that the maximization of information content within probability distributions can be well characterized by entropy, serving as a fundamental tool for quantifying uncertainty and disorder across various systems and distributions [14, 15]. In particular, the above findings are shown here to be highly advantageous for improving forecasting potential of elemental GBM process, by adopting entropy corrections to uncover patterns and structures within data sets of interest. The crux of introduced methodology lies in manipulating entropy measure through the strategic addition of trajectories, simulated by the traditional GBM method, to the existing time series distribution. By doing so, it is possible to notably refine our ability to predict and anticipate future data behavior.

### 16 Mar 2024 v2

# Entropy corrected geometric Brownian motion

@@ I Introduction
The geometric Brownian motion (GBM) has long served as a foundational model for capturing stochastic nature of systems characterized by the continuous random fluctuations. In particular, this model has has been argued to be well suited for forecasting diffusion processes Grebenkov et al. (2021), population dynamics Stojkoski et al. (2019) or most notably stock prices Black and Scholes (1973). However, central to GBM is the assumption that logarithm of its solutions result in a normal distribution, a premise that, while convenient for analytical purposes, may not fully encapsulate the complexities inherent in realistic systems Marathe and Ryan (2005). For instance, real data distributions often exhibit characteristics such as non-zero skewness, excess kurtosis or fluctuating volatility, notably deviating from the idealized bell curve Stojkoski et al. (2020). These deviations have significant implications for the accuracy of traditional GBM, by challenging its abilities to properly portray extreme events or interpret underlying dynamics of considered data.
By recognizing these limitations, a motivation arises for the exploration of solutions that go beyond the described constraints of normality. While alternative models, such as the stochastic volatility Heston (1993) and the jump-diffusion process Merton (1976), have been proposed, they mostly call for the modification or replacement of GBM. Hence, the existing advantages of this framework, in terms of its simplicity, efficiency and continuity, are somewhat hindered or lost. As an alternative, the present study proposes non-violent approach of employing entropy constraints to improve predictive capabilities of GBM without sacrificing its beneficial features. The idea originates from the Shannon’s information theory, where entropy plays pivotal role and stands for the minimum number of logical states needed to communicate some message Shannon (1948). This is to say, the well-ordered message is expected to have a lower entropy than a more random one. As a result, it can be argued that entropy provides a means to judge how well GBM predicts future events by measuring the level of uncertainty around forecast data with respect to the original time series. This premise is additionally reinforced when invoking the ability of entropy measure to capture extreme events Drzazga-Szczȩśniak et al. (2023).
To better picture above reasoning let us conceive gedanken experiment that hinges on a dice roll. We begin with a well known fact that probability of getting any side number during a conventional dice roll is equally probable, resulting in a uniform distribution. According to Shannon, the information about such process can be quantified by the entropy given as Shannon (1948):
@@ II Methodology
where $S(t)$ is the series value at time $t$, $\mu$ and $\sigma$ are respectively the drift and volatility coefficients, while $dW(t)$ denotes the Wiener process (the Brownian motion). By using Itô’s calculus and assuming log-normal distribution of considered data under GBM, the central equation of $S(t)$ can be obtained Itô (1944, 1951a, 1951b):
@@ III Biased dice roll
As a measure of the difference between the reference and generated probability distributions, given by GBM and EC-GBM methods, the Kullback-Leibler ($D_{KL}$) divergence is calculated and shown in Figure 3. The $D_{KL}$ measure is often used in information theory and statistics to quantify the difference between two probability distributions in the following manner Kullback and Leibler (1951):
@@ V Conclusion
In summary, the present research endeavor allows to deepen current understanding of how to effectively discern and harness the wealth of information contained within historical time series. It aligns with the concept that the maximization of information content within probability distributions can be well characterized by entropy, serving as a fundamental tool for quantifying uncertainty and disorder across various systems and distributions Jaynes (1957); Gupta et al. (2021). In particular, the above findings are shown here to be highly advantageous for improving forecasting potential of elemental GBM process, by adopting entropy corrections to uncover patterns and structures within data sets of interest. The crux of introduced methodology lies in manipulating entropy measure through the strategic addition of trajectories, simulated by the traditional GBM method, to the existing time series distribution. By doing so, it is possible to notably refine our ability to predict and anticipate future data behavior.